# Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models

### v1

@@ Appendix D Fitness Prompt for Judge-LLM
Appendix E Evaluation Prompt
Evaluation Prompt Template Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction: {input} ### Response:
Figure 6: Prompt template for code evaluation on HumanEval and MBPP
Appendix F Distribution of the Tasks
Figure 7: Task distribution of the generated instructions on roughly 600K samples of Genetic-Instruct out of 900K. Considering the domain of seed data, we select keywords that correspond to coding challenges. It is not an exhaustive list of the tasks and they can be overlapping.
Appendix G Length Distribution
Figure 8: Length distribution of the generated instructions, and code outputs on 900K samples of Genetic-Instruct. Instruction lengths are clipped to 300 words, while code outputs are clipped at 500 words for better visualization.
Appendix H Genetic-Instruct Algorithm
Genetic-Instruct Algorithm {python} import numpy as np task_ids = 0: "cross-over", 1: "mutation" task_prob = [0.5, 0.5] # 50 NUM_SAMPLES, NUM_COLONIES, NUM_SUBCOLONY = 500_000, 20, 10 def genetic_instruct(seed_instructions: List[Dict[str, str]]): """ Genetic Instruct for one generation, run on multiple nodes """ max_samples = NUM_SAMPLES / NUM_COLONIES current_samples = [] all_samples = [v for v in seed_instructions] while len(current_samples) <= max_samples: # Parallelize across NUM_SUBCOLONY processes (sub-colonies) futures = parallelize(genetic_instruct_one_iter, (all_samples,) for _in range(NUM_SUBCOLONY) )) results = await_all(futures) # wait for all sub-colonies to finish for result in results: current_samples.extend(result) all_samples.extend(result) # Deduplicate across all colonies current_samples = global_deduplicate_across_nodes(current_samples) write_dataset(current_samples) def genetic_instruct_one_iter(seed_instructions: List[Dict[str, str]]): """ One iteration of the Genetic Instruct algorithm Inputs: seed_instructions: List of (instruction, code) pairs from seed data + previous generations. Outputs: List of (instruction, code) pairs. """ # (1) Task Selection task_id = np.random.choice(task_ids.keys(), p=task_prob) # (2) Instruction Generation - sample a batch of seed data prompts = encode_inst_generation_prompts(seed_instructions, task_id) new_instructions = llm.generate(prompts) # (3) Code Generation - generate code for new inst, filter generations code_prompts = encode_codegen_prompts(new_instructions, task_id) code_generations = llm.generate(code_prompts) new_inst_code = filter_code_generations(new_instructions, code_generations) # (4) Code Quality Judgement - judge inst-code pairs for correctness judge_prompts = encode_judge_prompts(new_inst_code) code_judgements = llm.generate(judge_prompts) final_samples = [new_inst for(new_inst, judgement) in zip(new_instructions, code_judgements) if judgement == ’yes’] return final_samples
Figure 9: Algorithm of genetic-instruct as pseudo-code

### v2

@@ Appendix D Fitness Prompt for Judge-LLM
Appendix E Decontamination Prompt
Prompt Template for Contamination Detection Help me determine if the following two coding problems are the same. First problem: {instruction 1} Second problem: {instruction 2} Disregard the names and minor changes in word order that appear within. If the two problems are very similar and if they produce the same answer, we consider them to be the same problem. Respond with only "True" (problems are the same) or "False" (problems are different). Do not respond with anything else.
Figure 7: Prompt template for checking contamination
Appendix F Evaluation Prompts
Evaluation Prompt Template for MBPP and MBPP+ Here is a problem for which you need to generate code: {instruction} Please continue to complete the code with python programming language. The solution should be in the following format: “‘python # Your code here “‘ Do not generate any tests. Your function should have the same name as the function in the assert statement.
Figure 8: Prompt template for code evaluation on MBPP and MBPP+
Evaluation Prompt Template for HumanEval and HumanEval+ Here is a problem for which you need to complete code: {instruction} Please continue to complete the code with python programming language. The solution should be in the following format: “‘python # Your code here “‘ Do not generate any tests. You are not allowed to modify the given code and do the completion only.
Figure 9: Prompt template for code evaluation on HumanEval and HumanEval+